OpenStudy (anonymous):

A jet plane lands with a speed of 107.1 m/s and accelerates at a rate of −5 m/s^2 as it comes to rest. From the instant the plane touches the runway, what is the time needed before it can come to rest? Answer in units of s

OpenStudy (anonymous):

\[\large {v_i=107.1\ m/s \\ v_f = 0\ m/s\\ a = -5\ m/s^2 }\] Look through your formulas for something that has \[\huge v_i, v_f, v_a, \ {\large \textrm{and}} \ \Delta t\]

OpenStudy (anonymous):

Va=(Vf+Vi)/2

OpenStudy (anonymous):

Sorry, I meant \(\large a\), not \(\large v_a\).

OpenStudy (anonymous):

So, look for a formula that has all of the following: \(\huge v_i, v_f, a, \ {\large \textrm{and}} \ \Delta t\)

OpenStudy (anonymous):

a= Vf - Vi / Δt

OpenStudy (anonymous):

Yup. Now just plug in your values and solve for \(\Delta t\).

OpenStudy (anonymous):

okay:) thank you:)